OpenStudy (anonymous):

solve please 2/3x-2=5/6

OpenStudy (pandora):

2/3 becomes 4/6 4/6x -2 =5/6. Add 2 to both sides 4/6 x = 35/6. Then divide and you get x = 8.75

OpenStudy (anonymous):

how did it become 4/6?

OpenStudy (pandora):

You need common denominators so both must be 6. To make the denominator of 2/3 six you have to multiply the bottom by two. Thus you must also multiply the top by two making it the equivalent fraction of 4/6

OpenStudy (anonymous):

oh find the lcd?

OpenStudy (pandora):

Yes sorry I just said it the more complicated way

OpenStudy (anonymous):

my answer book says x=17/4